A company's net income is 100. Its sales are 1000, and assets and equity at the beginning of the year are 400 and 200, respectively. If it pays half its earnings as dividend, 1- what is the ROE at the end of the year? 2- what is the average asset turnover for the year 3- what is the leverage ratio at the end of year. 4- what is the return on equity if half of the net income of the year is used to pay down debt, and no dividends are paid.
Answer #1

1) ROE at the end of year

Shareholder's equity will change from 200 to 250 as 50% of the income shall be retained.

ROE = Net income / Shareholder's equity

ROE = 100/250

ROE =40%

2) Average asset turnover

=Net sales / Average total assets

=1000/400

=2.5

3) Leverage ratio at the end of year

Shareholders equity + Debt = Total assets

250+Debt=400

Debt=150

Debt to equity = 150/250 = 0.6

4) ROE if half of the net income of the year is used to pay down debt, and no dividends are paid:

Income = 100

Debt payment =100/2 = 50

Net income =100 - 50 = 50

ROE = Net income / Shareholder's equity

ROE = 50/200

ROE=25%